# Ratnangshu Das Robert Bosch Centre for Cyber-Physical Systems IISc, Bengaluru, India

### v1

@@ 4 Controller Synthesis
This section presents the proposed approximation-free, closed-form control design procedure.
Figure 1: Robot inside circular STT
@@
$e_{d}(t,x)$ $=\frac{\|x(t)-\mathsf{c}(t)\|}{\mathsf{r}(t)},\quad e_{\theta}(t,\xi)=\Psi\left(\frac{e_{d}(t,x)}{\overline{e}_{d}}\right)\frac{2}{\pi}\left(\tan^{-1}\left(\frac{\mathsf{c}_{2}(t)-x_{2}(t)}{\mathsf{c}_{1}(t)-x_{1}(t)}\right)-\theta\right).$
@@
Given the funnel functions, we then define normalized error $\hat{e}=[\hat{e}_{d},\hat{e}_{\theta}]^{\top}$ and transformed error $\varepsilon=[\varepsilon_{d},\varepsilon_{\theta}]^{\top}$ as:
$\hat{e}_{d}(t,\xi)=\frac{e_{d}(t,x)}{\rho_{d}(t)},\quad\hat{e}_{\theta}(t,\xi)=\frac{e_{\theta}(t,\xi)}{\rho_{\theta}(t)},\quad\varepsilon_{d}(t,x)=\log\left(\frac{1+\hat{e}_{d}(t,\xi)}{1-\hat{e}_{d}(t,\xi)}\right),\quad\varepsilon_{\theta}(t,\xi)=\log\left(\frac{1+\hat{e}_{\theta}(t,\xi)}{1-\hat{e}_{\theta}(t,\xi)}\right).$
@@
$v(t,\xi)$ $=-k_{d}\left(\varepsilon_{d}\alpha_{d}\cos(\psi-\theta)+\varepsilon_{\theta}\alpha_{\theta}\sin(\psi-\theta)\right),$
@@
$\dot{e}_{d}\!=\!\frac{(x_{1}-\mathsf{c}_{1})(v\cos\theta+d_{1}-\dot{\mathsf{c}}_{1})+(x_{2}-\mathsf{c}_{2})(v\sin\theta+d_{2}-\dot{\mathsf{c}}_{2})}{e_{d}\mathsf{r}^{2}}\!-\!\frac{e_{d}\dot{\mathsf{r}}}{\mathsf{r}},\dot{e}_{\theta}\!=\!\frac{2}{\pi}\Big(\frac{(x_{1}-\mathsf{c}_{1})(v\sin\theta+d_{2}-\dot{\mathsf{c}}_{2})-(x_{2}-\mathsf{c}_{2})(v\cos\theta+d_{1}-\dot{\mathsf{c}}_{1})}{e_{d}^{2}\mathsf{r}^{2}}\!-\!(\omega\!+\!d_{\theta})\Big).$ We also define the constraints for $\hat{e}$ through the open and bounded set $\mathbb{D}:=(-1,1)^{2}$.
@@
$\dot{V}$ $=\varepsilon_{d}\dot{\varepsilon}_{d}+\varepsilon_{\theta}\dot{\varepsilon}_{\theta}=(\varepsilon_{d}\alpha_{d}\cos(\psi-\theta)+\varepsilon_{\theta}\alpha_{\theta}\sin(\psi-\theta))v-(\varepsilon_{\theta}\alpha_{\theta}e_{d}\mathsf{r})\omega+\varepsilon_{d}\alpha_{d}\phi_{d}+\varepsilon_{\theta}\alpha_{\theta}\phi_{\theta},$
where $\phi_{d}=\frac{(x_{1}-\mathsf{c}_{1})(d_{1}-\dot{\mathsf{c}}_{1})+(x_{2}-\mathsf{c}_{2})(d_{2}-\dot{\mathsf{c}}_{2})}{e_{d}\mathsf{r}^{2}}-\frac{e_{d}\dot{\mathsf{r}}}{\mathsf{r}}-\hat{e}_{d}\dot{\rho}_{d}$ and $\phi_{\theta}=\frac{(x_{1}-\mathsf{c}_{1})(d_{2}-\dot{\mathsf{c}}_{2})+(x_{2}-\mathsf{c}_{2})(d_{1}-\dot{\mathsf{c}}_{1})}{e_{d}^{2}\mathsf{r}}-d_{\theta}-\frac{\pi}{2}\hat{e}_{\theta}\dot{\rho}_{\theta}$.
@@
$\dot{V}$ $=-k_{d}(\varepsilon_{d}\alpha_{d}\cos(\psi-\theta)+\varepsilon_{\theta}\alpha_{\theta}\sin(\psi-\theta))^{2}-k_{\theta}(\varepsilon_{\theta}\alpha_{\theta})^{2}e_{d}\mathsf{r}+\varepsilon_{d}\alpha_{d}\phi_{d}+\varepsilon_{\theta}\alpha_{\theta}\phi_{\theta}.$
Let, $\zeta=[\cos(\psi-\theta)\alpha_{d},\ \sin(\psi-\theta)\alpha_{\theta}]^{\top}$ and $\phi=[\phi_{d},\ \phi_{\theta}]^{\top}$. Then $\dot{V}\leq-k_{d}\|\varepsilon\|^{2}\|\zeta\|^{2}-k_{\theta}\|\varepsilon_{\theta}\alpha_{\theta}\|^{2}+\|\varepsilon\|\|\zeta\|\|\phi\|$. Adding and subtracting $k_{d}\Theta\|\varepsilon\|^{2}\|\zeta\|^{2}$ for some $\Theta\in(0,1)$:
$\dot{V}$ $\leq-k_{d}(1-\Theta)\|\varepsilon\|^{2}\|\zeta\|^{2}-k_{\theta}\|\varepsilon_{\theta}\alpha_{\theta}\|^{2}-\|\varepsilon\|\|\zeta\|\left(k_{d}\Theta\|\varepsilon\|\|\zeta\|-\|\phi\|\right).$
Therefore, $\dot{V}<0$ whenever
$\|\varepsilon\|>\frac{\|\phi\|}{k_{d}\Theta\|\zeta\|}.$
From Step 1, it holds that $\hat{e}_{d},\hat{e}_{\theta}\in(-1,1)$, implying the boundedness of $\xi(t)$ by bounds independent of $\tau_{\max}$. Therefore, since $d(\xi,t)$ is locally Lipschitz in $\xi$ and uniformly bounded in $t$, it is bounded for all $t\in[0,\tau_{\max})$. Moreover, since the tube center $\mathsf{c}(t)$, the tube radius $\mathsf{r}(t)$, and the funnel signals $\rho_{d}(t)$ and $\rho_{\theta}(t)$ are uniformly bounded, with $\mathsf{r}(t)>0$ for all $t$, we conclude that the remainder term $\phi=[\phi_{d},\phi_{\theta}]^{\top}$ is uniformly bounded by a finite constant $C_{\phi}$ (independent of $\tau_{\max}$) for all $t\in[0,\tau_{\max})$. Further, since $\hat{e}_{d},\hat{e}_{\theta}\in(-1,1)$ and $e_{d}\geq\bar{e}_{d}$, the corresponding gains $\alpha_{d}$ and $\alpha_{\theta}$ are uniformly lower bounded by positive constants. By further invoking $\zeta=[\cos(\psi-\theta)\alpha_{d},\ \sin(\psi-\theta)\alpha_{\theta}]^{\top}$, we conclude that $\|\zeta(t)\|\geq\zeta_{\min}>0$ for all $t\in[0,\tau_{\max})$. Hence, the threshold $\|\phi\|/(k_{d}\Theta\|\zeta\|)$ is bounded above by $\varepsilon^{*}=C_{\phi}/(k_{d}\Theta\zeta_{\min})$, implying the existence of a time-independent upper bound $\varepsilon^{*}\in\mathbb{R}_{0}^{+}$ on the transformed error $\varepsilon$, such that $\|\varepsilon(t)\|\leq\varepsilon^{*},\quad\forall t\in[0,\tau_{\max}).$
@@
$-1<\frac{e_{i}^{-\varepsilon^{*}}-1}{e_{i}^{-\varepsilon^{*}}+1}=:e_{i,L}\leq e_{i}\leq e_{i,U}:=\frac{e_{i}^{\varepsilon^{*}}-1}{e_{i}^{\varepsilon^{*}}+1}<1,$
$\forall t\in[0,\tau_{\max}),\ \text{for }i\in\{d,\theta\}.$ Therefore, by employing the control law (4.1), we can constrain $e$ to a compact subset of $\mathbb{D}$ as: $e(t)\in[e_{L},e_{U}]=:\mathbb{D}^{\prime}\subset\mathbb{D},\forall t\in[0,\tau_{\max})$, where $e_{L}=[e_{d,L},e_{\theta,L}]^{\top}$ and $e_{U}=[e_{d,U},e_{\theta,U}]^{\top}$.
@@
We know that $e(t)\in\mathbb{D}^{\prime},\forall t\in[0,\tau_{\max})$, where $\mathbb{D}^{\prime}$ is a non-empty compact subset of $\mathbb{D}$. However, if $\tau_{\max}<\infty$ then according to [sontag, Proposition C.3.6], $\exists t^{\prime}\in[0,\tau_{\max})$ such that $e(t)\notin\mathbb{D}$. This leads to a contradiction! Hence, we conclude that $\tau_{\max}$ can be extended to $\infty$.
@@
The time-varying control law in (4.1) provides a closed-form and approximation-free solution that guarantees fulfillment of the T-RAS task, even in the presence of unknown disturbances in control-affine systems. Moreover, the proposed algorithm can be extended to account for high-order dynamics by following the backstepping-like approach outlined in [verginis2022kdf].
@@ 5 Case Studies
To demonstrate the effectiveness of the proposed approach, we present two case studies.
@@ 5.1 Navigation in a Cluttered Office-Like Environment
In the first scenario, the robot navigates through a cluttered office-like environment with multiple static obstacles. The task is defined as a sequence of T-RAS objectives with given start, target, and unsafe regions, as shown in Fig. 2(d).
The synthesized spatiotemporal tube (STT) is shown in Fig. 2(a), and the corresponding robot trajectory under the proposed control law is shown in Fig. 2(d). The distance and orientation errors, $e_{d}$ and $e_{\theta}$, along with their funnel bounds $\rho_{d}(t)$ and $\rho_{\theta}(t)$, are plotted in Fig. 2(b)–(c). Both errors remain strictly within their funnels, confirming that the robot safely and successfully completes the sequence of reach-avoid-stay tasks within the prescribed time.
5.2 Reach-Avoid Task with Time-Varying Unsafe Set
The second scenario involves a dynamic environment, where the robot must reach the target region within a prescribed time while avoiding both static and moving obstacles. Figure 3(a) shows the synthesized STT, and Figs. 3(b)–(d) illustrate the workspace at three different time instants. The results demonstrate that the proposed controller effectively adapts to a changing environment by reshaping the tube, enabling the robot to avoid time-varying obstacles and reach the target within the specified time.
Figure 4: Comparison with existing approaches
Comparison: We compare the proposed STT-based control with state-of-the-art approaches such as CBF [toulkani2022reactive] and MPC [sun2017disturbance]. As shown in Fig. 4, all methods guide the robot from the same initial region to the target while avoiding obstacles. However, the proposed STT-based approach achieves this with significantly lower computation time (0.7 s) compared to CBF (3.3 s) and MPC (19.3 s). Moreover, under external disturbances, the CBF controller fails to prevent the robot from entering unsafe regions, whereas the proposed STT-based controller keeps the trajectory strictly within the tube, ensuring robust and safe navigation throughout the task.

### v2

@@ 4 Controller Synthesis
This section presents the proposed closed-form robust control design procedure.
Figure 1: Robot inside circular STT.
@@
$e_{d}(t,x)$ $=\frac{\|x(t)-\mathsf{c}(t)\|}{\mathsf{r}(t)}$ (10)
$e_{\theta}(t,\xi)$ $=\Psi\left(\frac{e_{d}(t,x)}{\overline{e}_{d}}\right)\frac{2}{\pi}\left(\tan^{-1}\left(\frac{\mathsf{c}_{2}(t)-x_{2}(t)}{\mathsf{c}_{1}(t)-x_{1}(t)}\right)-\theta\right).$
@@
Given the funnel functions, we then define normalized error $\hat{e}=[\hat{e}_{d},\hat{e}_{\theta}]^{\top}$ and transformed error $\varepsilon=[\varepsilon_{d},\varepsilon_{\theta}]^{\top}$ as follows:
$\hat{e}_{d}(t,\xi)$ $=\frac{e_{d}(t,x)}{\rho_{d}(t)},$ $\varepsilon_{d}(t,x)=\log\left(\frac{1+\hat{e}_{d}(t,\xi)}{1-\hat{e}_{d}(t,\xi)}\right),$
$\hat{e}_{\theta}(t,\xi)$ $=\frac{e_{\theta}(t,\xi)}{\rho_{\theta}(t)},$ $\varepsilon_{\theta}(t,\xi)=\log\left(\frac{1+\hat{e}_{\theta}(t,\xi)}{1-\hat{e}_{\theta}(t,\xi)}\right).$
@@
$v(t,\xi)$ $=k_{d}\left(\varepsilon_{d}\alpha_{d}\cos(\psi-\theta)-\varepsilon_{\theta}\alpha_{\theta}\sin(\psi-\theta)\right),$
@@
where $\dot{e}_{d}=\frac{(x_{1}-\mathsf{c}_{1})(v\cos\theta+d_{1}-\dot{\mathsf{c}}_{1})+(x_{2}-\mathsf{c}_{2})(v\sin\theta+d_{2}-\dot{\mathsf{c}}_{2})}{e_{d}\mathsf{r}^{2}}-\frac{e_{d}\dot{\mathsf{r}}}{\mathsf{r}},\\ \dot{e}_{\theta}=\frac{2}{\pi}\Big(\frac{(x_{1}-\mathsf{c}_{1})(v\sin\theta+d_{2}-\dot{\mathsf{c}}_{2})-(x_{2}-\mathsf{c}_{2})(v\cos\theta+d_{1}-\dot{\mathsf{c}}_{1})}{e_{d}^{2}\mathsf{r}^{2}}-(\omega+d_{\theta})\Big).$ We also define the constraints for $\hat{e}$ through the open and bounded set $\mathbb{D}:=(-1,1)^{2}$.
@@
$\dot{V}$ $=\varepsilon_{d}\dot{\varepsilon}_{d}+\varepsilon_{\theta}\dot{\varepsilon}_{\theta}$
$=-(\varepsilon_{d}\alpha_{d}\cos(\psi-\theta)-\varepsilon_{\theta}\alpha_{\theta}\sin(\psi-\theta))v-(\varepsilon_{\theta}\alpha_{\theta}e_{d}\mathsf{r})\omega+\varepsilon_{d}\alpha_{d}\phi_{d}+\varepsilon_{\theta}\alpha_{\theta}\phi_{\theta},$
where $\phi_{d}=\frac{(x_{1}-\mathsf{c}_{1})(d_{1}-\dot{\mathsf{c}}_{1})+(x_{2}-\mathsf{c}_{2})(d_{2}-\dot{\mathsf{c}}_{2})}{e_{d}\mathsf{r}}-e_{d}\dot{\mathsf{r}}-\mathsf{r}\hat{e}_{d}\dot{\rho}_{d}$ and $\phi_{\theta}=\frac{(x_{1}-\mathsf{c}_{1})(d_{2}-\dot{\mathsf{c}}_{2})-(x_{2}-\mathsf{c}_{2})(d_{1}-\dot{\mathsf{c}}_{1})}{e_{d}\mathsf{r}}-e_{d}\mathsf{r}d_{\theta}-\frac{\pi}{2}e_{d}\mathsf{r}\hat{e}_{\theta}\dot{\rho}_{\theta}$.
@@
$\dot{V}$ $=-k_{d}(\varepsilon_{d}\alpha_{d}\cos(\psi-\theta)-\varepsilon_{\theta}\alpha_{\theta}\sin(\psi-\theta))^{2}-k_{\theta}(\varepsilon_{\theta}\alpha_{\theta})^{2}e_{d}\mathsf{r}+\varepsilon_{d}\alpha_{d}\phi_{d}+\varepsilon_{\theta}\alpha_{\theta}\phi_{\theta}.$
To establish a strict upper bound on $\dot{V}$, we utilize the algebraic inequality $(A-B)^{2}\leq-\frac{1}{2}A^{2}+B^{2}$. Applying this to the linear velocity control term yields:
$\dot{V}$ $\leq-\frac{1}{2}k_{d}\varepsilon_{d}^{2}\alpha_{d}^{2}\cos^{2}(\psi-\theta)+k_{d}\varepsilon_{\theta}^{2}\alpha_{\theta}^{2}\sin^{2}(\psi-\theta)-k_{\theta}\varepsilon_{\theta}^{2}\alpha_{\theta}^{2}e_{d}\mathsf{r}+|\varepsilon_{d}|\alpha_{d}|\phi_{d}|+|\varepsilon_{\theta}|\alpha_{\theta}|\phi_{\theta}|$
$=-\frac{1}{2}k_{d}\cos^{2}(\psi-\theta)\alpha_{d}^{2}\varepsilon_{d}^{2}-\big(k_{\theta}e_{d}\mathsf{r}-k_{d}\sin^{2}(\psi-\theta)\big)\alpha_{\theta}^{2}\varepsilon_{\theta}^{2}+|\varepsilon_{d}|\alpha_{d}|\phi_{d}|+|\varepsilon_{\theta}|\alpha_{\theta}|\phi_{\theta}|.$
A crucial property of the proposed orientation funnel is that $\hat{e}_{\theta}\in(-1,1)$ guarantees $e_{\theta}\in(-\rho_{\theta,0},\rho_{\theta,0})$. From the definition of $e_{\theta}$, this implies $|\psi-\theta|\leq\frac{\pi}{2}\rho_{\theta,0}$. Since the initial funnel width $\rho_{\theta,0}<1$, the angular error is bounded away from $\pm\pi/2$. Thus, $\cos^{2}(\psi-\theta)\geq\cos^{2}(\frac{\pi}{2}\rho_{\theta,0}):=c_{\min}>0$.
Furthermore, since we are analyzing Case 2 where $e_{d}\geq\bar{e}_{d}$, we have $e_{d}\mathsf{r}\geq\bar{e}_{d}\mathsf{r}_{\min}$. By selecting the orientation control gain such that $k_{\theta}>k_{d}/(\bar{e}_{d}\mathsf{r}_{\min})$, we ensure the coefficient of $\varepsilon_{\theta}^{2}$ is strictly positive. Let $\lambda_{d}=\frac{1}{2}k_{d}c_{\min}$ and $\lambda_{\theta}=k_{\theta}\bar{e}_{d}\mathsf{r}_{\min}-k_{d}>0$. The derivative of the Lyapunov function is bounded by:
$\dot{V}\leq-\lambda_{d}\alpha_{d}^{2}\varepsilon_{d}^{2}-\lambda_{\theta}\alpha_{\theta}^{2}\varepsilon_{\theta}^{2}+|\varepsilon_{d}\alpha_{d}||\phi_{d}|+|\varepsilon_{\theta}\alpha_{\theta}||\phi_{\theta}|.$
Applying Young’s Inequality ($|xy|\leq\frac{\Theta}{2}x^{2}+\frac{1}{2\Theta}y^{2}$) to the remainder terms with $\Theta=\lambda_{d}$ and $\Theta=\lambda_{\theta}$ respectively:
$\dot{V}$ $\leq-\frac{\lambda_{d}}{2}\alpha_{d}^{2}\varepsilon_{d}^{2}-\frac{\lambda_{\theta}}{2}\alpha_{\theta}^{2}\varepsilon_{\theta}^{2}+\frac{\phi_{d}^{2}}{2\lambda_{d}}+\frac{\phi_{\theta}^{2}}{2\lambda_{\theta}}.$
From Step 1, since $\hat{e}_{d},\hat{e}_{\theta}\in(-1,1)$, the state $\xi(t)$ is constrained by bounds independent of $\tau_{\max}$. Since the tube parameters $\mathsf{c}(t),\mathsf{r}(t)$, the funnel signals $\rho_{d}(t),\rho_{\theta}(t)$, and the disturbances $d(\xi,t)$ are uniformly bounded, the remainder terms $\phi_{d}$ and $\phi_{\theta}$ are uniformly bounded by a finite constant $C_{\phi}$. Let $D_{\phi}=C_{\phi}^{2}/(2\lambda_{d})+C_{\phi}^{2}/(2\lambda_{\theta})$.
Since $\alpha_{d}$ and $\alpha_{\theta}$ contain the terms $(1-\hat{e}_{d}^{2})^{-1}$ and $(1-\hat{e}_{\theta}^{2})^{-1}$ respectively, they are bounded from below by positive constants $\underline{\alpha}_{d},\underline{\alpha}_{\theta}>0$. Therefore:
$\dot{V}\leq-\frac{\lambda_{d}\underline{\alpha}_{d}^{2}}{2}\varepsilon_{d}^{2}-\frac{\lambda_{\theta}\underline{\alpha}_{\theta}^{2}}{2}\varepsilon_{\theta}^{2}+D_{\phi}.$
This implies that $\dot{V}<0$ whenever $\|\varepsilon\|$ is sufficiently large. Specifically, there exists a time-independent upper bound $\varepsilon^{*}\in\mathbb{R}_{0}^{+}$ such that $\|\varepsilon(t)\|\leq\varepsilon^{*},\ \forall t\in[0,\tau_{\max})$.
@@
$-1<\frac{\exp(-\varepsilon^{*})-1}{\exp(-\varepsilon^{*})+1}=:e_{i,L}\leq\hat{e}_{i}\leq e_{i,U}:=\frac{\exp(\varepsilon^{*})-1}{\exp(\varepsilon^{*})+1}<1,$
$\forall t\in[0,\tau_{\max}),\ \text{for }i\in\{d,\theta\}.$ Therefore, by employing the control law (4.1), we can constrain $\hat{e}$ to a compact subset of $\mathbb{D}$ as: $\hat{e}(t)\in[\hat{e}_{L},\hat{e}_{U}]=:\mathbb{D}^{\prime}\subset\mathbb{D},\forall t\in[0,\tau_{\max})$, where $\hat{e}_{L}=[e_{d,L},e_{\theta,L}]^{\top}$ and $\hat{e}_{U}=[e_{d,U},e_{\theta,U}]^{\top}$.
@@
We know that $e(t)\in\mathbb{D}^{\prime},\forall t\in[0,\tau_{\max})$, where $\mathbb{D}^{\prime}$ is a non-empty compact subset of $\mathbb{D}$. However, if $\tau_{\max}<\infty$ then according to [21, Proposition C.3.6], $\exists t^{\prime}\in[0,\tau_{\max})$ such that $e(t)\notin\mathbb{D}$. This leads to a contradiction! Hence, we conclude that $\tau_{\max}$ can be extended to $\infty$.
@@
The time-varying control law in (4.1) provides a closed-form solution that guarantees fulfillment of the T-RAS task, even in the presence of unknown disturbances in control-affine systems. Moreover, the proposed algorithm can be extended to account for high-order dynamics by following the backstepping-like approach outlined in [26].
@@ 5 Case Studies
To demonstrate the effectiveness of the proposed approach, we present two case studies. All the computations were performed using Python 3.10 on a machine with a Windows operating system with an Intel Core i7-14700 CPU, 32 GB RAM. We have used the Z3 solver [8] to solve the optimization programs. The implementation of the proposed method can be found at: https://github.com/FocasLab/DifferentialDriveSTT.
@@ 5.1 Navigation in a Cluttered Office-Like Environment
In the first scenario, the robot has to navigate through a cluttered office-like environment with multiple static obstacles. The task is defined as a sequential reach-avoid objectives. For a $100$ second mission, the robot has to start within the ball ${\mathbf{S}}:=\mathcal{B}([3.5,0.8],0.2)$ and should reach its first target ${\mathbf{T}}_{1}:=\mathcal{B}([4,5.6],0.2)$ within $40$ seconds, while the next target ${\mathbf{T}}_{2}:=\mathcal{B}([8.4,3.6],0.2)$ in the next 30 seconds and finally reach its goal ${\mathbf{T}}_{3}:=\mathcal{B}([10,0.6],0.2)$ within the prescribed time while avoiding all the static obstacles in the environment. Given the sequential reach-avoid-stay tasks, we choose the center vector of STTs for individual tasks as two polynomials in time $t$ with the basis functions being the monomials of $t$ such as $\{1,t,t^{2},\ldots\}$ for two dimensions. We consider piecewise polynomial tubes and consider the degree of polynomials to be $4,6,4$ for the first, second and third mission respectively. Considering the obstacles to be static, $\mathcal{L}_{U}=0$. The algorithm converges with $\eta=-0.1$, with $\epsilon=0.25$. The obtained Lipschitz constant of the STT is $\mathcal{L}_{c}=0.2316,\mathcal{L}_{r}=0$. Therefore, $\eta^{*}+\mathcal{L}\epsilon=-0.1+0.2316\times 0.25=-0.0421<0$, satisfying the condition of Theorem 3.3, thereby implying the STT is formally verified.
[…]
5.2 Reach-Avoid Task with Dynamic Unsafe Set
[…]
Figure 4: Comparison with existing approaches.
Comparison: We compare the proposed STT-based control with state-of-the-art approaches such as CBF [25] and MPC [22]. As shown in Fig. 4, all methods guide the robot from the same initial region to the target while avoiding obstacles. However, the proposed STT-based approach achieves this with significantly lower per-step online control-synthesis time compared to CBF and MPC. Moreover, under external disturbances considered to be sinusoidal with the magnitude ranging between $0.0001$ to $0.0125$, the CBF and MPC based controller fails in some scenarios to prevent the robot from entering unsafe regions, whereas the proposed STT-based controller keeps the trajectory strictly within the tube, ensuring robust and safe navigation throughout the task. The average online computation time per step and success rate of the algorithms with and without disturbances are presented in the Table 1.
Table 1: Comparison Table of STT with classical algorithms
[…]
Without Disturbance With Disturbance Without Disturbance With Disturbance
STT $1\pm 0.0 1.1$$1\pm 0.0$ $100.00$ $100.00$
CBF [25] $395.34\pm 43.8$ $399.1\pm 19.2$ $100.00$ $90.00$
MPC [22] $63583.62\pm 1505.51$ $63500.64\pm 1146.84$ $100.00$ $44.00$